### v2

# Is Less More? Exploring Token Condensation as Training-free Adaptation for CLIP

Abstract
Contrastive language-image pre-training (CLIP) has shown remarkable generalization ability in image classification. However, CLIP sometimes encounters performance drops on downstream datasets during zero-shot inference. Test-time adaptation methods attempt to mitigate this by adjusting normalization layers, or tuning context prompts with large batch sizes and extensive augmentations, yet these methods are computationally intensive. This raises an important question: Is there a training-free approach that can efficiently address CLIP’s performance drop in such cases? To explore this, we benchmark token condensation techniques, originally designed to enhance the efficiency of vision transformers, on CLIP zero-shot inference tasks. We observe that although token condensation may compromise in-domain accuracy, it surprisingly enhances CLIP’s performance on certain cross-dataset benchmarks. This motivates two key inquiries: (1) Can token condensation serve as a “free-lunch” solution for CLIP zero-shot inference? (2) What criteria should guide condensation—how can essential tokens be identified and redundant ones eliminated? To address these questions, we propose Token Condensation as Adaptation (TCA), a training-free adaptation method for CLIP by pruning class-irrelevant visual tokens while merging class-ambiguous tokens. As the first approach for CLIP’s token efficiency, TCA demonstrates superior performance across cross-dataset tasks, achieving up to a 21.4% improvement over the strongest baseline while reducing GFLOPs by 12.2% to 48.9%, with minimized hyperparameter dependency. Our code is available at here.
1 Introduction
Online test-time adaptation (TTA) [54] has emerged as a promising strategy to handle distribution shifts encountered during inference [25]. TTA dynamically fine-tunes pretrained models on unlabeled data batches, enhancing generalization by aligning intermediate-layer batch statistics [34], optimizing for first-order flatness in the loss landscape [12], promoting self-supervised consistency across augmentations [66], or tracking model historical weights [24]. Despite the success of traditional TTA methods, they often require computationally expensive tuning of the backbone’s parameters. This challenge is further amplified in CLIP [36], which consist of vast parameter sets and require large batch sizes (e.g., 256) for stabilizing adaptation [8].
Test-time prompting (TPT) offers an efficient alternative for TTA by shifting adaptation focus to the language side of CLIP, learning a small set of task-specific context prompts only for downstream tasks. Nevertheless, TPT largely overlooks the impact of visual distribution shifts. Adapting to high-variance target images through prompts often relies on external source data [42] or extensive data augmentation [11] (e.g., 60$\times$ more AugMix or diffusion-based synthetic samples). In strict online TTA settings, where the batch size is constrained to one, this reliance on augmentation significantly inflates computational costs, leading to a 60$\times$ rise in GFLOPs compared to single-sample processing (i.e., $1108.61$ vs. $17.59$ GFLOPs). The need for gradient backpropagation during inference further increases the computation burden, making exiting TPT suboptimal for many resource-constrained applications.
What if we could address these challenges and adapt the CLIP model efficiently without additional training? In this paper, we take a closer look at whether training-free solutions can achieve this goal. Given that many CLIP visual encoders rely on Vision Transformers (ViTs), we explore ViT acceleration strategies such as token pruning and merging to reduce the number of redundant tokens as not all tokens contribute equally to predictions. While prior studies have often reported that reducing GFLOPs comes at the expense of significant performance degradation in in-distribution validation such as ImageNet-1K, our findings indicate that condensing tokens with lower attentiveness to its <cls> token can not only maintain but even improve CLIP’s performance on certain unseen target datasets (Tab. 1). To deepen our understanding, we conducted a preliminary leave-one-out analysis to investigate the role of individual visual tokens in CLIP’s visual-text alignment (Fig. 2(a)). The empirical results motivate us to target two specific types of tokens for condensation: (1) class-irrelevant background tokens, which may mislead CLIP by emphasizing non-essential regions that differ from the pre-training data; (2) class-ambiguous object tokens, such as cat whiskers or fur, which can overlap across other classes and disperse visual embeddings. However, relying solely on the current sample’s <cls> token for token condensation may not be optimal, as it tends to be overly universal and may not specifically align with the target classes. Although the text encoder in CLIP contains valuable target semantic information, the dimensional discrepancy between textual and visual embeddings restricts direct alignment. To find <cls> tokens that better represent target classes, we trace the historical <cls> tokens from target data streams on per class with the lowest uncertainty across time steps, which we refer to as domain <cls> tokens. As shown in Fig. 2(b), the maintained domain <cls> tokens follow a promising trajectory that gradually aligns with the text embeddings within the same visual space, bridging visual and textual representations, and guiding <cls> attentiveness to be more semantically aligned with target domains.
Building on these insights, we propose Token Condensation as Adaptation (TCA), a domain-aware token condensation strategy to adapt the CLIP model to unseen domains. As shown in Fig. 1, for every time step, the most reliable <cls> tokens are retained as domain <cls> tokens in a domain-aware token reservoir (DTR). These domain <cls> tokens, in turn, act as implicit indicators for token condensation. Specifically, for the current visual sample, we select inattentive tokens by considering both current <cls> tokens and domain <cls> tokens. Then, we utilize consensus across attention heads to retain only the most relevant class-specific patches and merge class-ambiguous tokens by coreset selection algorithm. Empirically, we show that a minimal number of centers is sufficient to stabilize performance, allowing TCA to scale efficiently with linear complexity. Besides, the domain <cls> tokens are explicitly used as token-level classifiers to correct CLIP logits and improve the final visual-text alignment. To our knowledge, this is the first work to explore token condensation in the context of test-time adaptation. The efficiency and effectiveness of the proposed training-free TCA have been rigorously validated across ten fine-grained classification datasets. In extensive evaluations with traditional TTA, prompting, and test-time prompting approaches, our method consistently outperforms the baselines by up to 21.4% compared with the strongest baseline, while reducing GFLOPs cost by 12.2% to 48.9%. For reproducibility, we provide source code in the supplementary material for reference.
Figure 1: An illustration of the proposed Token Condensation as Adaptation (TCA). To adapt CLIP visual embeddings to text embeddings during test-time, TCA utilizes a domain-aware token reservoir (DTR) to retain historical <cls> tokens with the lowest uncertainty as domain <cls>. These domain <cls> assist in (1) condensing tokens with low attentiveness scores (top-right) and (2) acting as token-level classifiers to refine predictions through logit correction, moving visual embeddings $\mathbf{z}_{t}$ toward text embeddings $\mathbf{t}_{c}$.
2 Related Work
Online Test-time Adaptation. To address performance degradation during test time, online test-time adaptation (TTA) has gained significant attention. Current TTA methods can be categorized into three main types [54]: optimization-based, data-based, and model-based approaches. Optimization-based methods focus on model updates and optimization objectives [14, 31, 66]. A prominent example is Tent [48], which adapts Batch Normalization layers [18] using entropy minimization. SAR [34] extends this approach to Layer Normalization [1] and Group Normalization [59] with sharpness-aware minimization [12]. Data-based methods include augmentations like selective augmentation [51] and adversarial augmentation [47], as well as the use of memory banks [13, 63, 5]. Model-based approaches involve architectural modifications to enhance model adaptability during testing [27, 19, 20, 53]. However, these methods typically depend on large batch sizes and augmentations, which introduce significant latency for online prediction.
Recently, vision-language models like CLIP [36] have excelled beyond fixed label sets, rendering traditional TTA methods less suitable [8]. As a result, various online adaptation strategies have been proposed to improve zero-shot generalization. Test-time prompt tuning has emerged as a key approach in this context. TPT [45] optimizes learnable prompts using data augmentations and soft entropy minimization, Diff-TPT [11] enriches this with more diverse augmentations [40], while C-TPT [62] focusing on model calibration. Other methods like VTE [8] and DART [28] leverage prompt ensembles with DART further employing moving averages to boost performance. SwapPrompt [29] incorporates an EMA-updated target prompt. AdaPrompt [65] utilizes a class-balanced memory bank to enhance adaptability. SCP [52] builds on TPT with a teacher-student framework to prevent semantic drift, while RLCF [67] incorporates reinforcement learning strategy [56] to optimize the adaptation process. Beyond these, MTA [64] introduces a new objective based on test-time augmentation to optimize visual features in the semantic space. TDA [21] further improves CLIP’s zero-shot ability by incorporating positive and negative caches with a training-free adapter. However, it relies on a large number of hyperparameters and is highly sensitive to them, while incurring significant computational costs during inference. In contrast, our approach strikes a better balance between computational efficiency and performance, outperforming both training-required and training-free methods.
Token Condensation in Vision Transformers. Vision transformers have achieved notable success in image recognition tasks, but their deployment is often limited by resource-constrained environments. To address this, various token condensation methods [32, 38, 41, 61, 70, 22, 50] have been proposed to reduce the computational overhead, primarily through two strategies: token pruning and token merging. Token pruning eliminates less informative tokens to save computation, as seen in methods like EViT [26], which retains tokens based on their attentiveness to the <cls> tokens. ATS [9] introduces input-dependent token pruning to adapt to variability across inputs. Token merging, on the other hand, seeks to combine similar tokens to reduce redundancy. For instance, ToME [3] uses bipartite soft matching to merge neighboring tokens that exhibit similarity. Hybrid approaches have also emerged, such as TPS [55], which prunes tokens and transfers information to retained ones using nearest-neighbor matching, and PruMerge [44], which prunes inattentive tokens using interquartile range and merges via k-nearest neighbors. While previous works have focused on enhancing efficiency within pure ViT models, our approach utilizes token condensation from a different perspective: addressing multimodal distribution shifts in VLMs. This shift remains underexplored, particularly in how to use semantic guidance to prune irrelevant visual tokens that introduce ambiguity. By condensing these tokens, we effectively reduce such distribution shifts, enhancing test-time performance while simultaneously lowering computational costs.
3 Our Approach
Problem Set-up. We begin by revisiting online test-time adaptation of CLIP. For a given downstream task $\mathcal{D}_{\operatorname{tar}}$, the test data $\mathbf{x}=\{\mathbf{x}_{t}\}_{t=1}^{T}$ arrives sequentially at each time step $t$. The objective is to adapt CLIP on the fly to classify the incoming test samples into one of $C$ classes, each represented by a textual prompt like “a photo of a <classname>”. CLIP embeds both visual and textual inputs into a shared space. The visual encoder $E_{v}$ extracts visual features $\mathbf{z}_{t}=E_{v}(\mathbf{V}_{t})\in\mathbb{R}^{D}$ from image patches $\mathbf{V}_{t}=[\mathbf{v}_{\operatorname{cls}},\mathbf{v}_{1},\dots,\mathbf{v}_{N}]\in\mathbb{R}^{(N+1)\times D_{v}}$ of dimension $D_{v}$, where $\mathbf{v}_{\operatorname{cls}}$ is a special <cls> token appended to $N$ patches. The text encoder $E_{t}$ generates class embeddings $\mathbf{T}=\{\mathbf{t}_{c}\}_{c=1}^{C}$, where each $\mathbf{t}_{c}\in\mathbb{R}^{D}$ corresponding to a class prompt. Classification is performed by computing the cosine similarity between the visual embedding $\mathbf{z}_{t}$ and each class embedding $\mathbf{t}_{c}$ with the probabilities calculated as:
$\mathbf{p}_{t,c}(\mathbf{z}_{t},\mathbf{t}_{c})=\frac{\exp\left(\operatorname{cos}(\mathbf{z}_{t},\mathbf{t}_{c})/\tau\right)}{\sum_{j=1}^{C}\exp\left(\operatorname{cos}(\mathbf{z}_{t},\mathbf{t}_{j})/\tau\right)},$ (1)
where $\tau$ denotes the temperature parameter controlling the sharpness of the output distribution.
3.1 Preliminary Studies
Pitsfalls of TTA. Since the target domain $\mathcal{D}_{\operatorname{tar}}$ is unseen during CLIP’s pre-training, the alignment between visual embeddings $\mathbf{z}_{t}$ and the textual embeddings $\mathbf{T}$ may be suboptimal. Previous methods have attempted to address this by learning domain-specific prompts [62] or replacing classifier weights with visual centroids [19] to move $\mathbf{T}$ closer to $\mathbf{z}_{t}$. However, the variability in CLIP’s visual embeddings is often much higher than in textual embeddings [36]. At the patch level, individual tokens within the visual embeddings can drift and vary significantly [36]. Thus, it becomes more urgent to derive methods that adjust $\mathbf{z}_{t}$ towards $\mathbf{T}$ for improved alignment.
(a) Impact of Token Removal on Alignment. The warmer color indicates higher attentiveness.
(b) Alignment between the updated domain <cls> token and text embedding over time.
Figure 2: Preliminary studies of token influence and the strategy of caching domain <cls> to improve alignment.
How to Mine <cls> Tokens Aligned with Text? In CLIP pretraining, the <cls> token is trained to align with a vast array of concepts, leading to attentiveness patterns that may extend beyond the target classes. Given the mismatch in dimensionality between textual and visual tokens, a fundamental challenge remains in finding a more representative <cls> token that consistently aligns with the text embeddings. To address this, we exploit temporal cues by tracing target samples with the lowest uncertainty as domain <cls>. The similarity between these domain <cls> tokens and $\mathbf{T}$ is plotted in Fig. 2(b). Our findings indicate that retaining these domain <cls> tokens with minimal uncertainty leads to a steady convergence of visual embeddings with the corresponding text embeddings in the same visual space. These domain <cls> tokens serve as contextually relevant proxies, effectively bridging the gap between visual and textual representations and enabling <cls> attentiveness to align more consistently with target classes.
3.2 Token Condensation as Adaptation
Building upon our empirical findings, we propose a novel strategy called Token Condensation as Adaptation (TCA), which selectively removes or merges tokens contributing to drift, enabling efficient adaptation in a training-free manner. Specifically, given an $L$-layer ViT, the forward pass through the $l$-th Transformer block, where $l\in[1,2,\dots,L]$, is formulated as:
$\mathbf{V}^{l+1}=\hat{\mathbf{V}}^{l}+\operatorname{MLP}(\hat{\mathbf{V}}^{l}),$ (2)
$\hat{\mathbf{V}}^{l}=\mathbf{V}^{l}+\frac{1}{H}\sum_{h=1}^{H}\operatorname{Attention}(\mathbf{V}^{l}\mathbf{W}_{Q}^{h},\mathbf{V}^{l}\mathbf{W}_{K}^{h})\mathbf{V}^{l}\mathbf{W}_{V}^{h},$
where $\mathbf{V}^{l}\in\mathbb{R}^{(N+1)\times D_{v}}$ represents the token embeddings at layer $l$. The matrices $\mathbf{W}^{h}_{Q}$, $\mathbf{W}^{h}_{K}$, $\mathbf{W}^{h}_{V}\in\mathbb{R}^{D_{v}\times D_{v}}$ represent linear projection for the query, key, value vectors in the $h$-th attention head, and $H$ denotes the total number of attention heads. To effectively condense tokens without compromising performance, we introduce two key functions: the token pruning function $f_{\operatorname{prune}}$ and the token merging function $f_{\operatorname{merge}}$. In line with [26], these functions are applied between the multi-head self-attention layers and feed-forward layers of each transformer block. The modified forward pass becomes:
$\hat{\mathbf{V}}^{l}=f_{\operatorname{merge}}\circ f_{\operatorname{prune}}\left(\mathbf{V}^{l};\mathfrak{R}\right),$ (3)
where $f_{\operatorname{prune}}(\cdot;\mathfrak{R}):\mathbb{R}^{N+1}\mapsto\mathbb{R}^{(\alpha\cdot R\cdot N)+1}$ and $f_{\operatorname{merge}}(\cdot;\mathfrak{R}):\mathbb{R}^{(\alpha\cdot R\cdot N)+1}\mapsto\mathbb{R}^{R\cdot N+1}$ are responsible for reducing the number of tokens from $N+1$ (including the <cls> token) to $R\cdot N+1$, where $R$ represents the fraction of tokens to be preserved. The parameter $\alpha$ controls the extent of token pruning, ensuring only the most semantically significant tokens are retained. Detailed explanation of $f_{\operatorname{prune}}$ and $f_{\operatorname{merge}}$ can be found in Sec. 3.3 and Sec. 3.4.
Domain-aware token reservoir $\mathfrak{R}$. Notably, both $f_{\operatorname{prune}}$ and $f_{\operatorname{merge}}$ rely on a class-specific token reservoir =$\mathfrak{R}=\{\mathfrak{R}_{c}\}_{c=1}^{C}$. Each buffer $\mathfrak{R}_{c}=\{(\mathbf{H}_{c}(\mathbf{z}_{i};\mathbf{t}_{c}),\mathbf{A}_{i,c}^{\operatorname{cls}})\}_{i=1}^{M}$ is structured as a priority queue that retains the top $M$ most reliable domain <cls> tokens of target samples, which serve to implicitly distil semantic information from the corresponding text prompt $\mathbf{t}_{c}$ to guide the visual adaptation. These domain <cls> tokens are crucial alignment proxies: although the architectures of the text encoder $E_{t}$ and the visual encoder $E_{v}$ differ, the selected domain <cls> tokens help determine which visual tokens best align with text features from CLIP’s perspective. The reliability of these domain <cls> tokens is quantified by entropy scores $\mathbf{H}_{c}(\mathbf{z}_{t},\mathbf{t}_{c})$,
$\mathbf{H}_{c}(\mathbf{z}_{t},\mathbf{t}_{c})=-\mathbf{p}_{t,c}\left(\mathbf{z}_{t},\mathbf{t}_{c}\right)\log\mathbf{p}_{t,c}\left(\mathbf{z}_{t},\mathbf{t}_{c}\right),$ (4)
which act as keys to update the reservoir $\mathfrak{R}_{c}$. At each time step $t$, for each visual embedding $\mathbf{z}_{t}$, the corresponding <cls> embeddings from all $L$ layers $\mathbf{A}_{t,c}^{\operatorname{cls}}=[\mathbf{v}_{\operatorname{cls}}^{1},\ldots,\mathbf{v}_{\operatorname{cls}}^{L}]\in\mathbb{R}^{L\times D_{v}}$ will be stored in $\mathfrak{R}_{c}$ if $\operatorname{argmax}(\mathbf{p}_{t,c})=c$, ensuring that only the most semantically consistent samples are retained:
$\mathfrak{R}_{c}\leftarrow\operatorname{update}\left(\mathfrak{R}_{c},\left(\mathbf{H}_{c}(\mathbf{z}_{t},\mathbf{t}_{c}),\mathbf{A}_{t,c}^{\operatorname{cls}}\right)\right).$ (5)
If the priority queue $\mathfrak{R}_{c}$ has reached its capacity $M$, the sample with the highest entropy score is discarded and replaced with the new sample. Strategies for updating the reservoir, such as first-in-first-out (FIFO) and similarity- or diversity-enforcing methods, are explored in Sec. 4.3.
Logits Self-correction. To counter the shifts on the semantic side, we introduce a logits self-correction mechanism that leverages domain <cls> tokens stored in $\mathfrak{R}$. In particular, the <cls> embedding of the current sample, $\mathbf{V}_{t}^{\operatorname{cls}}\in\mathbb{R}^{L\times D_{v}}$, is compared with the stored domain <cls> tokens denoted as a set $\mathcal{A}=\{\mathbf{A}_{i,c}^{\operatorname{cls}}\}_{i=1}^{M}$. The cosine similarity between these cross-layer <cls> tokens serves as a token-level classifier, which provides auxiliary information to adjust the predicted probability $\mathbf{p}_{t,c}$ from a visual perspective:
$\tilde{\mathbf{p}}_{t,c}=\mathbf{p}_{t,c}+\lambda\mathbf{p}^{\operatorname{token}}_{t,c},$ (6)
$\mathbf{p}^{\operatorname{token}}_{t,c}=\frac{1}{M}\sum_{i=1}^{M}\operatorname{cos}(\mathbf{V}_{t}^{\operatorname{cls}},\mathbf{A}_{i,c}^{\operatorname{cls}})\cdot\mathbf{P}\cdot\mathbbm{1}_{c},$
where $\lambda$ is the logit correction weight. $\mathbbm{1}_{c}\in\mathbb{R}^{C}$ the one-hot vector for the $c$-the class. The layer-specific exponential scaling coefficients are denoted as $\mathbf{P}=[\exp(\frac{l}{\beta})]_{l=1}^{L}\in\mathbb{R}^{L}$, where $\beta$ controls the influence of different layers. We show that this correction temperature $\beta$ provides semantic interpretability, as further discussed in Sec. 4.3. This self-correction mechanism ensures better alignment between the visual and semantic contexts, improving robustness in handling semantic shifts.
Figure 3: An overview of domain-aware cross-head token pruning.
3.3 Domain-aware Cross-head Token Pruning
Prior token pruning methods for ViTs such as [26] primarily discard patch tokens with lower averaged attention scores $\mathbf{S}\in\mathbb{R}^{N}$ relative to the <cls> token $\mathbf{v}_{\operatorname{cls}}^{l}$ across all attention heads. The pruning process can be expressed as
$\hat{\mathbf{V}}_{\operatorname{prune}}^{l}\leftarrow\{\hat{\mathbf{v}}_{i}^{l}~{}|~{}\mathbf{S}_{i}\leq\theta_{\operatorname{prune}}(\alpha,R),\forall i\in[N]\},$ (7)
$~{}\mathbf{S}_{i}=\frac{1}{H}\sum_{h=1}^{H}\operatorname{Attention}(\mathbf{v}_{\operatorname{cls}}^{l}\mathbf{W}_{Q}^{h},\mathbf{v}^{l}_{i}\mathbf{W}_{K}^{h}),$
where $\hat{\mathbf{V}}_{\operatorname{prune}}^{l}$ denotes the set of tokens retained after pruning at layer $l$. Here, the pruning threshold $\theta_{\operatorname{prune}}(\alpha,R)$ is determined by the desired pruning ratio, ensuring only top-ranked $\alpha\cdot R\cdot N$ tokens are retained. However, this approach faces two limitations when applied in TTA tasks: (1) The <cls> token is universal and may not be specifically aligned with the target class set. It may capture broad, unrelated semantics (e.g., “cat food”), leading to the retention of irrelevant tokens that mislead the model into making incorrect predictions of the target class (e.g., “cat”). (2) Averaging attention scores across all heads risks omitting important details, as each attention head tends to focus on distinct features (e.g., shape, color). Outliers in attention heads (highlighted by red circles in Fig. 3) may disproportionately dominate the overall score, overshadowing valuable information by other heads and leading to suboptimal pruning decisions. To overcome these limitations, we propose a cross-head token pruning function that evaluates the token importance individually for each attention head and utilizes the averaged relative ranking positions to determine which tokens to prune (see Fig. 3). This approach reaches a more robust cross-head consensus and mitigates the impact of outliers. To guide pruning with more representative <cls> tokens, we sample an domain <cls> token for the $(l-1)$-th layer from $\mathfrak{R}_{c^{*}}$, where $c^{*}$ is determined by comparing the cosine similarity between the current $\mathbf{v}^{l}_{\operatorname{cls}}$ token embedding and the stored domain <cls> tokens $\mathbf{A}^{l-1}_{i,c}\in\mathcal{A}$:
$\mathbf{A}_{c^{*}}^{l-1}=\frac{1}{M}\sum\nolimits_{i\in[M]}\mathbf{A}_{i,c^{*}}^{l-1},$ (8)
where $\mathbf{A}_{c^{*}}^{l-1}\in\mathbb{R}^{D_{v}}$ is the averaged domain <cls> token sampled from $\mathfrak{R}_{c^{*}}$, providing historical context that is better aligned with target semantics. $c^{*}=\arg\max\nolimits_{c\in[C]}\operatorname{cos}(\mathbf{v}_{\operatorname{cls}}^{l},\mathbf{A}_{i,c}^{l-1}),$. Subsequently, we refine the attention map by inserting this historical domain <cls> token $\mathbf{A}_{c^{*}}^{l-1}$ to compute the pruning scores $\mathbf{S}^{\operatorname{head}}_{i}$ for the $i$-th token:
$\mathbf{S}^{\operatorname{head}}_{i}=\frac{1}{H}\sum_{h=1}^{H}\operatorname{rank}_{h}(i),$ (9)
$~{}\hat{\mathbf{V}}_{\operatorname{prune}}^{l}\leftarrow\{\hat{\mathbf{v}}_{i}^{l}~{}|~{}\mathbf{S}^{\operatorname{head}}_{i}\leq\theta_{\operatorname{prune}}(\alpha,R),\forall i\in[N]\},$
where $\text{rank}_{h}(i)$ gives the relative ranking position of token $i$ in head $h$ based on the attention score to the concatenation of $\mathbf{v}^{l}_{\operatorname{cls}}$ and $\mathbf{A}_{c^{*}}^{l-1}$, computed as $\operatorname{Attention}([\mathbf{v}_{\operatorname{cls}}^{l};\mathbf{A}_{c^{*}}^{l-1}]\mathbf{W}^{h}_{Q},[\mathbf{V}^{l};\mathbf{A}_{c^{*}}^{l-1}]\mathbf{W}^{h}_{K})$. [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] indicates concatenation. This method ensures that tokens receiving consistently high attention across individual heads are retained, thereby achieving greater robustness to outliers in specific attention heads.
3.4 Domain-aware Token Merging
As depicted in Fig. 2(a), a subset of tokens, although relevant to the target class, exhibit high uncertainty. These tokens are referred to as class-ambiguous tokens, identified from the ranked token list derived using LABEL:eq:sort:
$\Phi=\{i~{}|~{}\theta_{\operatorname{merge}}(R)\leq\mathbf{S}^{\operatorname{head}}_{i}\leq\theta_{\operatorname{prune}}(\alpha,R),\forall i\},$ (10)
where $\theta_{\operatorname{merge}}(R)$ denotes thresholds for token selection during merging. The selected tokens $\mathbf{V}^{l}_{\Phi}=\{\mathbf{v}_{i}^{l}\}_{i\in\Phi}$ can introduce variance or noise into latent representation $\mathbf{z}_{t}$ and negatively impact the final classification decision. To address this, we propose a domain-aware token merging strategy to consolidate these tokens into more representative ones.
Existing token merging strategies either truncate neighbored token pairs with high similarity with bipartite soft matching [3] or apply spectral clustering [2], and graph pooling [58] to merge similar tokens at higher costs. In contrast, we adopt a more efficient coreset selection approach, which identifies the most representative tokens $\hat{\mathbf{V}}^{l}_{\operatorname{merge}}\in\mathbb{R}^{K\times D_{v}}$ from $\mathbf{V}^{l}_{\Phi}$ and assigns the remaining ambiguous tokens to these selected tokens. The coreset selection strategy is equivalent to solving the K-Center problem [57, 43]. The objective is to select $K$ center tokens such that the maximum distance between any token and its nearest center is minimized. Formally, the greedy search for optimization is defined as follows:
$\mathbf{C}^{*}={\arg\min}_{\mathbf{C}\subseteq\mathbf{V}^{l}_{\Phi},|\mathbf{C}|=K}\max_{\mathbf{v}_{i}^{l}\in\mathbf{V}^{l}_{\Phi}}\min_{\mathbf{v}_{c}^{l}\in\mathbf{C}}d(\mathbf{v}_{i}^{l},\mathbf{v}_{c}^{l}),$ (11)
where $\mathbf{C}^{*}\in\mathbb{R}^{K\times D_{v}}$ represents the set of selected center tokens, $K$ is the number of centers, and $d(\cdot,\cdot)$ is the distance metric between token $\mathbf{v}_{i}^{l}$ and center token $\mathbf{v}_{c}^{l}$. Once the center tokens $\mathbf{C}^{*}$ are selected, the remaining tokens are assigned to their nearest centers, and the ambiguous tokens are merged as:
$\hat{\mathbf{V}}_{\operatorname{merged}}^{l}=\frac{1}{|\mathcal{N}(k)|}\sum\nolimits_{\mathbf{v}_{i}^{l}\in\mathcal{N}(k)}\mathbf{v}_{i}^{l},$ (12)
where $\mathcal{N}(k)$ represents the set of tokens assigned to center $k$. The value of $K$ is kept small, with $K\ll N$, allowing our merging algorithm to operate with linear complexity. The final embedding is composed of the class token $\mathbf{v}^{l}_{\operatorname{cls}}$, the pruned tokens $\hat{\mathbf{V}}^{l}_{\operatorname{prune}}$ excluding those in the class-ambiguous set $\Phi$, and the merged tokens $\hat{\mathbf{V}}^{l}_{\operatorname{merge}}$. This combined embedding is then passed to the next layer in the ViT. The study of $k$ and the details of the algorithm are provided in the supplementary material.
4 Experiments
4.1 Experimental Setup
Datasets. Following prior works, we conducted the cross-dataset (CD) benchmark. The CD benchmark assesses the model’s performance on unseen classes across 10 datasets: Aircraft [30], Caltech101 [10], Cars [23], DTD [6], EuroSAT [15], Flower102 [33], Food101 [4], Pets [35], SUN397 [60], and UCF101 [46]. Details of other benchmark experiments are provided in the supplementary material.
Baselines. We compare TCA with existing approaches across four categories: (1) Prompt-tuning methods like CoOp [69] and CoCoOp [68], which require multi-epoch adaptation; (2) Conventional online test-time adaptation (TTA) methods such as Tent [48] and SAR [34]. Tent updates batch normalization layers, while SAR further incorporates sharpness-aware minimization for reliable model updates. Following [8], we reran these experiments with adjusted batch sizes to align with our settings; (3) Test-time prompting methods, including TPT [45], C-TPT [62], and Diff-TPT [11], as well as TTA methods for CLIP such as MTA [64] and TDA [21]; and (4) Token pruning and merging methods for ViTs, such as EViT [26], ToMe [3], and ATS [9]. As ATS is an adaptive token pruning method with no fixed budget, we constrain its computational cost by an upper bound to ensure fair comparison.
Implementation Details. We utilize the official CLIP prompts as text inputs. The batch size is set to 1 without data augmentations to mimic realistic deployment scenarios. All experiments are conducted using the pre-trained CLIP models, specifically using ViT-B/16 and ViT-L/14 architectures as the visual backbone. We set $K$ to 2. Notably, our method is training-free, which achieves rapid adaptation with minimized hyperparameter dependency. All experiments are performed on one NVIDIA RTX A6000 GPU.
4.2 Main Results
Table 1: Results on the cross-dataset benchmark using CLIP ViT-B/16, including the number of learnable parameters (L-Param.) for learning-based TTA methods. denotes the averaged GFLOPs across all datasets.
Method
Aug-free
Aircraft
Caltech101
Cars
DTD
EuroSAT
Flower102
Food101
Pets
SUN397
UCF101
Average
GFLOPs
L-Param.
CLIP 23.22 93.55 66.11 45.04 50.42 66.99 82.86 86.92 65.63 65.16 64.59 17.59 0
CoOp 18.47 93.70 64.51 41.92 46.39 68.71 85.30 89.14 64.15 66.55 63.88 17.59 2048
CoCoOp 22.29 93.79 64.90 45.45 39.23 70.85 83.97 90.46 66.89 68.44 64.63 17.59 34,816
Tent 8.97 93.39 62.69 39.78 20.85 61.23 83.70 87.76 65.30 66.93 59.06 17.59 40,960
SAR 21.09 91.85 61.15 44.68 46.19 63.54 81.43 87.95 59.74 65.58 62.32 17.59 31,744
TPT 24.78 94.16 66.87 47.75 42.44 68.98 84.67 87.79 65.50 68.04 65.10 1108.61 2048
Diff-TPT 25.60 92.49 67.01 47.00 43.13 70.10 87.23 88.22 65.74 62.67 65.47 - -
C-TPT 23.90 94.10 66.70 46.80 48.70 69.90 84.50 87.40 66.00 66.70 65.47 1108.61 2048
MTA 25.32 94.21 68.47 45.90 45.36 68.06 85.00 88.24 66.67 68.11 65.53 - -
TDA 23.91 94.24 67.28 47.40 58.00 71.42 86.14 88.63 67.62 70.66 67.53 17.59 0
EViTR=0.9 24.12 92.25 64.57 45.09 48.41 70.24 84.99 88.96 64.58 68.46 65.17 15.41 0
ToMER=0.9 24.66 92.49 63.10 44.92 48.64 69.22 85.04 87.90 64.22 68.62 64.88 15.31 0
ATSR=0.9 22.86 92.21 57.90 40.96 40.62 67.52 80.16 85.34 61.53 67.22 61.63 11.15 0
EViTR=0.7 23.31 91.20 58.44 43.32 43.26 67.11 79.70 85.77 61.41 66.69 62.02 11.62 0
ToMER=0.7 22.26 90.79 55.48 42.32 40.12 64.11 79.36 84.19 60.66 63.97 60.33 11.45 0
ATSR=0.7 17.28 85.40 33.65 36.52 27.79 52.62 55.97 72.94 48.82 56.44 48.74 8.76 0
TCAR=0.9 24.87 93.63 65.33 46.16 70.43 73.33 85.31 89.53 65.92 72.38 68.69 $\mathbf{15.45}_{\mathbf{{\color[rgb]{0,0,1}-12.2\%}}}$ 0
TCAR=0.7 23.19 92.13 58.15 44.50 61.63 69.79 79.99 85.99 61.89 67.38 64.46 $\mathbf{11.69}_{\mathbf{{\color[rgb]{0,0,1}-33.5\%}}}$ 0
Tab. 1 presents the results for fine-grained cross-dataset benchmark using the ViT-B/16 architecture. As observed in Fig. 2(a), the core idea behind TCA is that condensing inattentive tokens can effectively mitigate distribution shifts caused by visual-text misalignment. This concept is first validated by the improved performance of token pruning baselines over CLIP inference, where a condensed token set yields a $0.9\%$ increase in average accuracy when $R=0.9$. TCA further enhances its performance by dealing with visual-text misalignment, moving visual features toward historical domain <cls> tokens from DTR. As a result, TCA achieves an average accuracy of 68.69%, outperforming both train-required and training-free baselines without augmentation. Conventional TTA methods perform poorly on all datasets even with the requirement of fine-tuning a large amount of learnable parameters. In contrast, prompt-tuning methods, although requiring fewer learnable parameters, rely heavily on augmentation and struggle to effectively handle visual shifts. While TDA is a training-free method, it requires a large number of hyperparameters (a total of 10 for managing positive and negative caches) to achieve optimal performance. On the other hand, TCA uses significantly fewer hyperparameters and delivers a 1.72% improvement in average accuracy over TDA, with approximately 12.2% fewer GFLOPs. Further details on TDA combined with token condensation baselines, as well as the impact of the visual backbone (ViT-L/14) and OOD benchmarks, are presented in supplementary material.
(a) Impact of reservoir size.
(b) Impact of GFLOPs budget.
Figure 4: Impact of reservoir size and GFLOPs budgets.
4.3 Ablation Study
We conducted a comprehensive ablation study to evaluate TCA’s effectiveness and efficiency. For further analysis on merging center and pruning to merging ratio, see the supplementary material.
Table 2: Impact of correction weight $\lambda$.
$\lambda$ 2 3 4 5 6 7 8
Pets 89.53 89.32 89.13 88.96 88.96 88.66 88.44
Flower102 72.43 72.76 73.20 73.16 73.16 73.33 73.16
EuroSAT 60.15 65.74 68.80 69.51 69.84 70.16 70.43
Table 3: Impact of scale factor $\beta$.
$\beta$ 0.01 0.05 1 3 5
Pets 89.51 89.53 89.37 89.42 89.26
Flower102 73.33 73.08 70.93 70.56 70.44
EuroSAT 63.64 64.06 69.86 70.26 70.43
Impact of Logits Correction Temperature $\beta$. Tab. 3 examines how different logit correction temperatures $\beta$ affect the adaptation results. The intuition is that with a smaller $\beta$ value, the logits correction will emphasize the tokens in shallower layers (Eq. 6), while a larger $\beta$ value will shift the focus to deeper layers. We observe that a smaller value of $\beta$ is preferred for the Pets dataset as it contains animals as objects, requiring more high-level contextual information for accurate predictions [37]. In contrast, for EuroSAT, the best predictions are obtained with larger $\beta$ values, suggesting that low-level, local information is crucial. This aligns well with the nature of the dataset, where different types of land can be distinguished by features such as colors and edges. Nevertheless, our method consistently provides significant improvements across all $\beta$ values, with accuracy gains of up to 20%, highlighting the effectiveness of logit correction using the domain <cls> tokens.
Impact of Correction Weight $\lambda$. To investigate how different correction weights $\lambda$ affect performance, as described in Eq. 6, we conducted experiments across a wide range of $\lambda$ values, from 2 to 8, as shown in Tab. 2. We observe that Pets exhibits stable results across different $\lambda$ values, indicating that less aggressive correction is sufficient. In contrast, datasets such as Flower102 and EuroSAT which initially do not perform well on CLIP, benefit from stronger corrections, achieving their best performance with larger correction weights of 7 and 8, respectively. This highlights the effectiveness of our logits correction module.
Impact of GFLOPs Budget. We evaluate TCA’s performance under different GFLOPs budgets: $R=\{0.6,0.7,0.8,0.9\}$, resulting in GFLOPs of 9.91, 11.68, 13.27, and 15.45, respectively, compared to the baseline ($R=1$, 17.58 GFLOPs). As shown in Fig. 4(b), condensing inattentive tokens can even enhance performance on certain datasets, notably Pets, and EuroSAT. Specifically for EuroSAT, when $R=0.9$, the model’s adaptation performance is significantly improved, aligning with our findings in Fig. 2(a). However, excessively aggressive pruning budgets (e.g., GFLOPs less than 13) lead to significant performance degradation across all datasets. This occurs since higher pruning rates may inadvertently remove informative tokens, causing irreversible harm in training-free scenarios where we lack supervision or the ability to update the model for extra correction.
Table 4: Impact of reservoir updating strategy.
FIFO Uncertainty Similarity-enf Diversity-enf
Pets 89.21 89.18 88.91 89.53
Flower102 70.65 73.28 72.15 73.33
EuroSAT 50.28 70.20 68.48 70.43
Impact of Reservoir Saving Strategy. In Tab. 4, we examine the performance changes across different reservoir saving strategies. We compare several approaches: First-In-First-Out (FIFO); an uncertainty-based strategy, which discards the most uncertain sample when the reservoir reaches capacity; a similarity-enforced strategy, where samples with high certainty and high cosine similarity to the saved samples are preferred; and a diversity-enforced strategy, which prioritizes saving prototypes that contain distinct tokens compared to those already stored. Our results show that the FIFO strategy performs poorly on Flower102 and EuroSAT, likely because CLIP’s low confidence leads to retaining misclassified samples. Conversely, Pets has high CLIP zero-shot accuracy (86.91% in Tab. 1), which makes FIFO acceptable. Among all strategies, the diversity-based approach consistently achieves the best performance. This is intuitive, as it maintains a representative set of features by capturing dataset diversity, whereas entropy-based methods may store homogenous features and overlook multiple class prototypes. By prioritizing diversity, our method ensures that a more representative set of features is maintained, leading to more robust performance across datasets.
Impact of Reservoir Size $M$. We assess the effectiveness of TCA across various reservoir sizes $M$ on Pets, Flower102, and EuroSAT datasets, as illustrated in Fig. 4(a). Remarkably, although the best performances are achieved at different reservoir sizes for different datasets, our TCA consistently maintains stable and high performance across a wide range of $M$ values. This showcases the robustness and flexibility of TCA with respect to different reservoir budgets. Notably, even under extreme conditions with a minimal reservoir size (i.e., $M=1$), our strategy significantly surpasses the strongest baseline method, TDA, by a large proportion on the EuroSAT dataset (14.2%).
Table 5: Impact of the proposed components in TCA.
$\mathbf{A}_{c^{*}}^{l}$ $\mathbf{S}^{\text{head}}$ Food101 Pets EuroSAT
85.25 88.77 68.14
85.31 88.96 67.69
85.25 89.23 67.14
85.31 89.53 70.43
Figure 5: Examples of our token condensation. More visualizations can be found in the supplementary material.
Impact of Component. The impact of the historical domain <cls> $\mathbf{A}_{c^{*}}^{l}$ and the head-wise sorting score $\mathbf{S}^{\text{head}}$ (Eq. 8) is presented in Tab. 5. We observe that each component individually contributes to performance improvements. On the Food101 and Pets datasets, incorporating either component yields measurable gains in accuracy. By leveraging historical domain <cls> tokens, the model acquires rich contextual information, enhancing the stability of token importance over time. Simultaneously, cross-head token sorting ensures that token pruning decisions are more robust by accounting for consensus across attention heads. An intriguing case arises with the EuroSAT dataset. Here, the baseline performance without any components is 68.14%. Applying either component alone results in a slight performance decrease. However, when both components are used together, performance significantly improves to 70.43%. This outcome emphasizes the necessity of combining historical domain <cls> tokens and cross-head token sorting to fully realize the model’s potential.
Visualization of Token Condensation. We visualize the pruned and merged tokens of different ViT layers in Fig. 5. The black mask indicates pruned regions while different colors are set for different merging clusters. As token condensation progresses, non-discriminative tokens are gradually removed, leading to better alignment with the text semantics. See the supplementary material for more details.
5 Conclusion
In this paper, we introduced Token Condensation as Adaptation (TCA), a novel training-free test-time adaptation method for CLIP models. Our comprehensive experiments demonstrated that token condensation significantly benefits the visual-text alignment in CLIP, which can further serve as an interpretation of visual semantics. Additionally, our method reduces GFLOPs as a beneficial byproduct, enhancing computational efficiency. For fair comparisons, we fixed the GFLOPs budget by pre-setting $R$ in our experiments; however, the condensing rate can be adaptively estimated using the distances to the merging center as an indicator, which can be a promising direction for future research. We also acknowledge the limitations of TCA as a training-free method, particularly its bottleneck in handling datasets with severe distribution shifts (refer to the supplementary material).
Supplementary Material
This supplementary material provides additional descriptions of the proposed TCA, including empirical results and algorithm. Visual aids for token condensation are also included to enhance understanding of the proposed method.
  • Sec. 5.1: Additional Experiments and Ablation Study.
  • Sec. 5.2: Token Condensation Algorithm.
  • Sec. 5.3: Quantitative Study ($R=0.7$).
  • Sec. 5.4: Potential Limitation of TCA.
5.1 Additional Results
Impact of Visual Backbone.
Table 6: Results on the cross-dataset benchmark with CLIP ViT-L/14. denotes the averaged GFLOPs across all datasets.
Method Aircraft Caltech101 Cars DTD EuroSAT Flower102 Food101 Pets SUN397 UCF101 Average GFLOPs
CLIP 31.59 94.56 78.12 57.03 63.00 79.58 90.92 93.46 69.05 76.13 73.34 81.14
Tent 27.45 94.97 76.93 57.15 66.20 74.83 89.20 93.27 68.73 75.73 72.45 81.14
SAR 26.07 94.52 75.58 56.91 63.77 75.03 89.13 93.05 68.39 75.50 71.80 81.14
TPT 30.06 95.21 76.84 52.30 55.11 76.21 88.56 93.08 67.69 73.78 70.88 143.31
TDA 33.42 95.46 78.72 57.39 66.27 79.94 90.83 93.27 70.74 78.14 74.42 81.14
EViTR=0.9 31.23 94.56 76.59 56.38 63.04 79.13 90.08 93.32 68.54 76.40 72.93 65.19
ToMER=0.9 28.29 92.54 71.26 56.68 60.30 77.87 89.77 91.28 68.21 72.22 70.84 64.74
ATSR=0.9 25.74 93.39 67.69 55.02 52.81 76.78 86.48 91.50 66.26 72.56 68.82 43.62
EViTR=0.7 26.94 92.94 62.55 53.96 52.04 73.24 80.69 90.00 63.70 71.21 66.73 40.78
ToMER=0.7 15.60 83.73 38.43 49.82 44.51 59.36 72.65 77.73 58.32 50.99 55.11 40.05
ATSR=0.7 6.87 67.87 16.37 40.78 30.12 37.43 34.50 60.94 30.07 33.44 35.84 26.76
TCAR=0.9 33.84 96.39 76.93 56.38 67.74 80.71 90.21 93.54 70.02 78.24 74.40 $\mathbf{65.24}_{\mathbf{{\color[rgb]{0,0,1}-19.6\%}}}$
TCAR=0.7 29.73 94.81 63.72 53.72 60.69 76.00 81.55 90.02 65.61 73.14 68.90 $\mathbf{41.44}_{\mathbf{{\color[rgb]{0,0,1}-48.9\%}}}$
Trends similar to ViT-B/16 are observed with the ViT-L/14 architecture, as shown in Tab. 6. TCA consistently surpasses TDA across multiple datasets, including Aircraft, Caltech101, EuroSAT, Flower102, Pets, and UCF101, while adhering to a limited GFLOPs budget (19.6% GFLOPs reduction). Even with a 48.9% reduction in GFLOPs, TCA continues delivering satisfactory results. This demonstrates the scalability and robustness of our method across different model sizes, reinforcing its effectiveness without additional training.
Table 7: Impact of token merging/pruning ratio.
Merging:Pruning 0:1 1:2 2:1
Pets 89.04 88.99 89.53
EuroSAT 69.63 69.98 70.43
Impact of Pruning & Merging Ratio. We experiment with different token pruning and merging ratios under the same computational budget, as shown in Tab. 7. Incorporating token diversity through merging consistently enhances performance. Specifically, the 2:1 merging-to-pruning ratio outperforms other configurations, especially those favoring pruning. This is because merging preserves diverse token representations by K coresets that pure pruning might discard. When comparing pruning-only (0:1) with the 1:2 merging-pruning ratio on Pets, pruning-only performs better. This may be because the dataset features images with a single prominent object, meaning that pruning background tokens has minimal impact since essential object information remains intact. In contrast, for the EuroSAT dataset, which comprises diverse satellite imagery, simply pruning tokens leads to the loss of important contextual features necessary for accurate classification.
Table 8: Impact of the merging center number K.
K 1 2 3 4
Pets 89.29 89.53 89.29 89.21
EuroSAT 66.25 70.43 66.96 67.44
Food101 85.15 85.31 85.31 85.38
Impact of Merging Center Number $K$. We evaluate TCA performance by giving different numbers of merging centers $K$ for Pets, EuroSAT, and Food101 datasets. As shown in Tab. 8, setting $K=2$ consistently yields the best results. This choice balances preserving important information and reducing redundancy. A smaller $K$ (i.e., $K=1$) may oversimplify the merging process, leading to the loss of critical details, especially in diverse datasets like EuroSAT. Conversely, increasing $K$ beyond 2 introduces unnecessary complexity and can over-segment the token space, retaining redundant tokens that contribute little to classification. Therefore, maintaining a very small $K$ (where $K\ll N$) is sufficient and advantageous.
Figure 6: Accuracy gain of TCA and TDA + EViT/ToME over TDA combined with ATS.
Comparison with State-of-the-Art TTA Using Token Condensation. We additionally evaluate the performance when combining TDA with token pruning and merging baselines and show the performance gain over TDA $+$ ATS in Fig. 6. Although TDA achieves considerable performance gain, it heavily relies on the negative cache and a large set of hyperparameters. In contrast, TCA’s accuracy gain significantly surpasses that of TDA $+$ EViT and TDA $+$ ToME across multiple datasets and on average, even with a minimal set of hyperparameters, highlighting its superior adaptation capability.
Table 9: Results on the out-of-distribution benchmark with CLIP ViT-B/16. denotes the averaged GFLOPs across all datasets.
Method Aug-free ImageNet ImageNet-A ImageNet-V2 ImageNet-R ImageNet-S Average OOD Average GFLOPs
CLIP $\checkmark$ 68.34 49.89 61.88 77.65 48.24 61.20 59.42 17.59
Tent 65.49 44.57 59.26 78.72 22.52 54.11 51.27 17.59
SAR 58.52 33.71 53.95 76.08 39.24 52.30 50.75 17.59
TPT 68.98 54.77 63.45 77.06 47.94 62.44 60.81 1108.61
Diff-TPT 70.30 55.68 65.10 75.00 46.80 62.28 60.52 -
C-TPT 69.30 52.90 63.40 78.00 48.50 62.42 60.70 1108.61
MTA 70.08 58.06 64.24 78.33 49.61 64.06 62.56 -
TDA 69.26 50.82 62.23 77.93 50.26 62.10 60.31 17.59
EViTR=0.95 68.32 49.46 61.73 77.00 47.76 60.85 58.99 16.31
ToMER=0.95 67.57 48.81 60.88 75.78 47.05 60.02 58.13 16.21
ATSR=0.95 65.83 49.80 59.47 71.09 43.38 57.91 55.94 11.50
TCAR=0.95 68.88 50.13 62.10 77.11 48.95 61.43 59.57 16.55
Impact of Benchmark Datasets. We conducted experiments on the OOD benchmark which focuses on evaluating the model’s effectiveness on shifted data using label sets previously seen by CLIP. This includes variants of ImageNet [7]: ImageNet-A [16], ImageNet-V2 [39], ImageNet-R [17], and ImageNet-S [49]. A consistent observation can be seen in the out-of-distribution (OOD) benchmark, where TCA demonstrates significant improvements over the CLIP baseline under a constrained GFLOPs budget of $R=0.95$, as shown in Tab. 9. TCA outperforms traditional test-time adaptation methods while maintaining efficiency. TCA also achieves superior results on ImageNet-R and ImageNet-S, outperforming TPT without augmentation. Additionally, when compared to other training-based approaches, even those with unlimited computational budgets, TCA delivers comparable performance. However, we observe that TCA does not perform as strongly on the OOD benchmark as it does on the CD benchmark even with a higher rate $R$. This may be due to the conceptual shifts in OOD datasets, as shown in Sec. 5.4, which could present a challenge for training-free adaptation methods.
5.2 Algorithm
Algorithm 1 Token Condensation at the $l$-Layer in $E_{v}$
1:
2:Token reservoir $\mathfrak{R}$;
3:Visual patches $\mathbf{V}^{l-1}$ at layer $l-1$;
4:Pruning threshold $\theta_{\text{prune}}(\alpha\cdot R)$;
5:Merging threshold $\theta_{\text{merge}}(R)$
6:Token-efficient visual feature $\hat{\mathbf{V}}^{l}$
7:Domain <cls> Token Selection: Obtain $\mathbf{A}^{l-1}_{c^{*}}$ by Eq. 8, using domain <cls> tokens in $\mathbf{\mathfrak{R}}$ and sample’s <cls> token $\mathbf{v}^{l}_{\text{cls}}$
8:Compute cross-head scores $\mathbf{S}_{i}^{\text{head}}$ for every token $i$
9:if  $\forall i$, $S_{i}^{\text{head}}\leq\theta_{\text{prune}}(\alpha\cdot R)$  then
10:       Token Pruning: Obtain $\mathbf{\hat{V}}^{l}_{\text{prune}}$ via LABEL:eq:sort
11:end if
12:if  $\forall i$, $\theta_{\text{merge}}(R)\leq S_{i}^{\text{head}}\leq\theta_{\text{prune}}(\alpha\cdot R)$  then
13:       Token Merging: Obtain $\mathbf{\hat{V}}^{l}_{\text{merged}}$ via Eq. 12
14:end if
15:return $\mathbf{\hat{V}}^{l}$, which is composed of $\mathbf{v}^{l}_{\operatorname{cls}}$, $\mathbf{\hat{V}}^{l}_{\text{prune}}$ (excluding merged tokens), and $\mathbf{\hat{V}}^{l}_{\text{merged}}$
Figure 7: Sample data from the OOD benchmark. The samples from the same class exhibit significant diversity. For instance, in the ImageNet-R dataset, one image of a great white shark is dominated by shoes and human legs, while another is on top of a building, showing extreme variability.
Algorithm 1 outlines the process for performing token pruning and merging at layer $l$ in a ViT-based CLIP model. We first obtain the averaged domain <cls> tokens $\mathbf{A}^{l-1}_{c^{*}}$ by the <cls> tokens saved in the reservoir $\mathfrak{R}$. Token condensation is then conducted given the domain <cls> token. Specifically, we conduct token pruning by relative ranking positions of token $i$ across multiple attention heads. Then, coreset selection is used for token merging. Finally, we concatenate the <cls> token $\mathbf{v}^{l}_{\operatorname{cls}}$ with the retained tokens as the input for the next layer, where the original $N+1$ tokens are shrunk to $(R\cdot N)+1$, thereby reducing the computational cost.
5.3 Quantitative Study
Figure 8: Visualization of our proposed token condensation with $R=0.7$. Pruned tokens are masked in black, while different colors represent distinct merging clusters.
We visualize the token condensation masks at layer 3, layer 6, and layer 9, and compare them with the original image across multiple datasets, as shown in Fig. 8. As the layers go deeper, we observe that class-irrelevant patches are gradually pruned, as indicated by the black mask. TCA also merges class-ambiguous patches, such as fur in cat images, and ground and sky in aircraft and car images. All similar tokens are merged into a single token using our proposed coreset selection strategy. After token condensation, the sample features retain only discriminative information, thereby bridging the gap between visual and text features, and mitigating the distribution shift between pretrained data and unseen datasets.
5.4 Discussion on the Limitation of TCA
In this section, we discuss the potential limitations of our proposed TCA. Due to the training-free nature of the approach, it is challenging to mitigate the performance gap when the testing domain diverges significantly from the training domain. As observed in the out-of-distribution (OOD) samples shown in Fig. 7, the ground truth object is not always centrally located, and larger class-irrelevant objects (e.g., humans or shoes) can sometimes dominate the prediction. This issue is particularly prominent in CLIP models, where text features for all classes are predefined. When the dominant object is included in the label set, accurately directing visual features to the correct class without additional training becomes difficult. Moreover, the diversity of OOD samples introduces further complexity, especially in the absence of data augmentation. These observations raise important questions for future research: (1) How can we quantify the capacity to mitigate domain shift effectively? (2) What lightweight solutions can be developed for backpropagation and network updates to facilitate test-time adaptation? We leave these questions for future work.